# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 1. Introduction
Lately, the Bessel periods on the symplectic group $\mathsf{G}:={\bf GSp}_{2}$ of rank $2$ progressively gain its importance in the arithmetic of Siegel modular forms. The notion of Bessel period and the associated local model of admissible representations of $\mathsf{G}$ were originally introduced by Novodvorski and Piatetski-Shaprio in their work ([26], [27], [30]), where a modern treatment of Andrianov’s integral representation ([1], [2]) of the spinor $L$-function for Siegel modular forms was developed. Related to this, we should also mention independent work by Sugano ([43]), which handles vector-valued Hilbert-Siegel cusp forms on arithmetic groups defined by maximal lattices of indefinite division algebras. The Bessel period also plays an essential role in a formulation of Böcherer’s conjecture posed by Liu [23] in a style of the refined Gan-Gross-Prasad conjecture. This version of Böcherer’s conjecture has been completely solved by Furusawa and Morimoto ([10], [11]).
In this paper, we compute the Rankin-Selberg type integral a la Piatetski-Shapiro for holomorphic vector-valued Siegel cusp forms of square-free levels explicitly, using results by Pitale-Schmidt [32] on local new vectors on $\mathsf{G}({\mathbb{Q}}_{p})\,(p<\infty)$; since we rely on the local-global method, our result is immediately applied to non-holomorphic Siegel cusp forms of discrete series type (cf. [20], [21]). We explain the moral of our investigation for scalar-valued Siegel modular forms of weight $l$. Our focus is an asymptotic formula as $N+l\rightarrow\infty$ of the average
@@
where $\mathsf{G}^{\#}$ is a subgroup of $\mathsf{G}$ defined by a quadratic field $E$ of discriminant $D<0$ (see §2.1), $E(s,\Lambda,\mu)$ is the Eisenstein series on $\mathsf{G}^{\#}({\mathbb{A}})$ attached to an ideal class character $\Lambda$ of $E$, a character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}^{\times}\rightarrow{\mathbb{C}}^{1}$ and $s\in{\mathbb{C}}$ (see §3.1), $B^{\Lambda}(\varphi)$ is the Bessel period of $\varphi$ in an orthonormal basis ${\mathscr{B}}(l,N)$ of Siegel cusp forms of weight $l$ and square free level $\Gamma_{0}(N)$, and $\langle E(s,\Lambda,\mu)\mid\varphi\rangle$ is the $L^{2}$-inner product of $E(s,\Lambda,\mu)$ and $\varphi|\mathsf{G}^{\#}({\mathbb{A}})$. The conductor of $\mu$ is assumed to be relatively prime to $2DN$. Actually, for application ([19]), we consider a more general average than (1.1) by requiring ${\mathscr{B}}(l,N)$ to consist of joint eigenforms by Hecke operators outside $NM$ and by putting the eigenvalue of such a Hecke operator at each $\varphi$ in the formation of the sum (1.1). With such a generality, we obtain a completely explicit expression of (1.1) in terms of
@@
  • An average of Fourier coefficient of a newform $\varphi_{\pi}^{0}$ in $\pi$, denoted by $R(\varphi_{\pi}^{0},E,\Lambda)$ (see (4.12)) introduced in [17]).
@@
For the precise statement including the vector-valued case, we refer to Proposition 5.1 and Theorem 5.2. Owing to the explicit nature of our result, we are able to show that, asymptotically as the prime level $N$ grows with the weight $l$ being fixed, the contribution to (1.1) of the old forms, the Yoshida lifts and the Saito-Kurokawa lifts are 00 (see Theorem 5.4). These results are necessary in our forthcoming paper [19], where we shall work out the geometric side of a relative trace formula whose spectral side is essentially (1.1). We should remark that an average similar to our computed average (5.9) without the factors ${\widehat{L}}(s+1/2,\pi,\mu)$ and $\widehat{f_{S}}(\pi)$ is studied in [7, §3]. When $E={\mathbb{Q}}(i)$, our average (5.9) is almost the same as the one in [5, §8] except the appearance of the local periods.
@@
Let us explain the structure of this paper. In §2, we introduce basic objects related to algebraic subgroups of the symplectic group $\mathsf{G}$ and Haar measures on their $p$-adic, real and adelic points. In §3.1, the Rankin-Selberg type integral and the global Bessel function for Siegel cusp forms are recalled; after this preparation, the basic identity (Lemma 3.2) is stated, which computes the pairing $\langle E(s,\Lambda,\mu),\varphi\rangle$ in (1.1) in terms of the Mellin transform of the global Bessel function attached to $\varphi$. We review the proof ([30]) to determine a constant depending on Haar measures exactly. In §4, we recall the definition of the quantity $R(\varphi,E,\Lambda)$ and describe its behavior under the Galois conjugation $\Lambda\mapsto\Lambda^{\dagger}$ (Lemma 4.6). The local multiplicity one theorem for Bessel models of $\pi_{p}$’s allows us to split $B^{\Lambda}(\varphi)$ for $\varphi\in\pi$ corresponding to a pure tensor in $\otimes_{p}\pi_{p}$ to an Euler product of local Bessel functions for $\pi_{p}$ up to a constant. To define our global newform $\varphi_{\pi}^{0}$ and identify this constant with $R(\varphi_{\pi}^{0},E,\Lambda)$ (see Lemma 4.7), we fix pairs of a local Bessel functional $\ell_{p}$ on $\pi_{p}$ and a local new vector $\xi_{p}$ of $\pi_{p}$ normalized by $\ell_{p}(\xi_{p})=1$; in §4.4, we explain that the main result of [32] ensures the existence of such $\{(\ell_{p},\xi_{p})\}_{p}$ when $\pi$ corresponds to a new forms on $\Gamma_{0}(N)$. In §6, we compute the local zeta-integrals by using results of [32]. The unramified computations are known as recalled in §6.2. Other cases, as well as the explicit determination of the local periods of prime level (see §6.8), seem to be new, so that some details are given in §6.5, §6.6, and §6.4. In §4.6, the functional equation of the spinor $L$-function is deduced from the basic identity by using Lemma 4.6. Although a “nice” functional equation of the spinor $L$-function itself is known in a broader generality ([40, Lemma 1.2]) owing to Arthur’s classification of automorphic representations of $\mathsf{G}({\mathbb{A}})$, we include this section for the sake of completeness and to provide a proof free from Arthur’s result, confirming that the local $L$-factor and the local $\varepsilon$-factor defined by [30] (in terms of a particular Bessel model) coincides with the ones proposed in [37, Tables 2,3]. In section §5, we prove Proposition 5.1 and Theorem 5.2; by invoking a result by Furusawa-Morimoto ([10, Theorem 8.1]), we obtain an explicit expression of (1.1) solely in terms of $L$-functions (Theorem 5.3).
@@ 2.1. Embedding of groups
Let $E={\mathbb{Q}}(\sqrt{D})$ be an imaginary quadratic field of discriminant $D<0$; we realize $E$ as a subfield of ${\mathbb{C}}$ so that $\sqrt{D}=i|D|^{1/2}$ with $i$ being the imaginary unit. Let $\mathfrak{o}_{E}$ denote the ring of integers of $E$. Choose a ${\mathbb{Z}}$-basis $\{1,\theta\}$ of $\mathfrak{o}_{E}$ so that
@@
The different of $E/{\mathbb{Q}}$, defined as ${\mathfrak{d}}_{E/{\mathbb{Q}}}^{-1}:=\{\tau\in E\mid{\operatorname{tr}}_{E/{\mathbb{Q}}}(\tau\mathfrak{o}_{E})\subset{\mathbb{Z}}\}$, is $\frac{1}{\sqrt{D}}\mathfrak{o}_{E}$; thus a symplectic ${\mathbb{Z}}$-bi-linear form $\langle\cdot,\cdot\rangle:\mathfrak{o}_{E}^{2}\times\mathfrak{o}_{E}^{2}\longrightarrow{\mathbb{Z}}$ is defined by
@@
Then $\iota_{\theta}(h)\in{\bf GL}_{4}(R)$ satisfies $\langle\iota_{\theta}(h)(x),\iota_{\theta}(h)(y)\rangle=\det(h)\,\langle x,y\rangle$ for all $x,y\in\mathfrak{o}_{E}^{2}$, i.e., $\iota_{\theta}(h)$ belongs to $\mathsf{G}(R)$ and $\nu(\iota_{\theta}(h))=\det h$. Thus, we have an embedding
@@
Let ${\mathsf{B}}^{\#}$ be the Borel subgroup of $\mathsf{G}^{\#}$ consisting of all those upper-triangular matrices in $\mathsf{G}^{\#}$; then, for any ${\mathbb{Q}}$-algebra $R$, the set ${\mathsf{B}}^{\#}(R)$ coincides with the set of all points of the form
@@ 2.2. Open compact subgroups at finite places
Let $p$ be a prime number. Set $E_{p}:=E\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ and $\mathfrak{o}_{E,p}:=\mathfrak{o}_{E}\otimes_{\mathbb{Z}}{\mathbb{Z}}_{p}$. Define
${\bf K}^{\#}_{p}:=\mathsf{G}^{\#}({\mathbb{Z}}_{p}),\quad{\bf K}_{p}:=\mathsf{G}({\mathbb{Z}}_{p}).$
Then, ${\bf K}^{\#}_{p}$, which coincides with the stabilizer in $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ of the $\mathfrak{o}_{E,p}$-lattice $\mathfrak{o}_{E,p}^{2}\subset E_{p}^{2}$, is a maximal compact subgroup of $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$, and ${\bf K}_{p}$ is the standard maximal compact subgroup of $\mathsf{G}({\mathbb{Q}}_{p})$. For a non-zero ideal ${\mathfrak{n}}\subset{\mathbb{Z}}_{p}$, set
@@
For a non-zero ideal ${\mathfrak{n}}\subset{\mathbb{Z}}_{p}$, we have
${\bf K}_{0}^{\#}({\mathfrak{n}}\mathfrak{o}_{E,p})=\iota_{\theta}^{-1}({\bf K}_{0}({\mathfrak{n}})).$
@@ 2.3. Maximal compact subgroup at the archimedean place
Note that ${\bf K}_{\infty}\subset{\bf Sp}_{2}({\mathbb{R}})$. The action of $\mathsf{G}({\mathbb{R}})^{0}$ on the Siegel upper-half space
${\mathfrak{h}}_{2}:=\{Z\in{\bf M}_{2}({\mathbb{C}})\mid{}^{t}Z=Z,\,{\operatorname{Im}}(Z)\gg 0\},$
denoted by $(g,Z)\mapsto g\langle Z\rangle$ is defined by the usual formula, i.e.,
@@ 2.4. Haar measures
For locally compact unimodular topological groups $H$ relevant to us, we fix Haar measures $\eta_{H}$ on $H$ in the following manner.
@@
On the Lie group $\mathsf{G}^{\#}({\mathbb{R}})=\{g^{\#}\in{\bf GL}_{2}({\mathbb{C}})\mid\det(g^{\#})\in{\mathbb{R}}^{\times}\}$, we fix a Haar measure as follows. By the Iwasawa decomposition, any element $g^{\#}\in\mathsf{G}^{\#}({\mathbb{R}})$ is uniquely written as
(2.19) $g^{\#}=\left[\begin{smallmatrix}1&\beta\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}a\tau&0\\ 0&\tau^{-1}\end{smallmatrix}\right]\,k^{\#}$
with $a\in{\mathbb{R}}^{\times}$, $\tau\in{\mathbb{C}}^{\times}$, $\beta\in{\mathbb{C}}$ and $k^{\#}\in{\bf K}_{\infty}^{\#}$. Then, our Haar measure on $\mathsf{G}^{\#}({\mathbb{R}})$ is symbolically defined as
(2.20) ${\rm{d}}g^{\#}=|a|_{\mathbb{R}}^{-2}|\tau|_{\mathbb{C}}^{-2}\times{\rm{d}}\eta_{{\mathbb{C}}}(\beta)\,\tfrac{{\rm{d}}\eta_{{\mathbb{R}}}(a)}{|a|_{\mathbb{R}}}\,\tfrac{{\rm{d}}\eta_{{\mathbb{C}}}(\tau)}{|\tau|_{\mathbb{C}}}\,\,{\rm{d}}k^{\#},$
where ${\rm{d}}k^{\#}$ is the Haar measure on ${\bf K}_{\infty}^{\#}$ with volume $1$, and $|\tau|_{\mathbb{C}}:=\tau\bar{\tau}$. For $p<\infty$, the group $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ is endowed with the Haar measure ${\rm{d}}g_{p}^{\#}$ uniquely determined by the condition ${\operatorname{vol}}({\bf K}_{p}^{\#})=1$. As in the archimedean case, if we write a general element $g^{\#}$ of $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ as in (2.19) with $a\in{\mathbb{Q}}_{p}^{\times}$, $\tau\in E_{p}^{\times}$, $\beta\in E_{p}$ and $k^{\#}\in{\bf K}_{p}^{\#}$, then
(2.21) ${\rm{d}}g^{\#}_{p}=|a|_{p}^{-2}|{\operatorname{N}}_{E_{p}/{\mathbb{Q}}_{p}}\tau|_{p}^{-2}\times{\rm{d}}\eta_{E_{p}}(\beta)\,{{\rm{d}}\eta_{{\mathbb{Q}}_{p}^{\times}}(a)}\,{{\rm{d}}\eta_{E_{p}^{\times}}(\tau)}\,{\rm{d}}k^{\#},$
where ${\rm{d}}k^{\#}$ is the normalized Haar measure of ${\bf K}^{\#}_{p}$. On the adele group $\mathsf{G}^{\#}({\mathbb{A}})$, we use the product measure of Haar measures on $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ for $p\leqslant\infty$. Then, from (2.20) and (2.21), we get
(2.22) $\int_{\mathsf{G}^{\#}({\mathbb{A}})}f(g^{\#})\,{\rm{d}}g^{\#}=\tfrac{\sqrt{|D|}}{2}\int_{\mathsf{N}^{\#}({\mathbb{A}})}\int_{{\mathbb{A}}^{\times}}\int_{{\mathbb{A}}^{\times}_{E}}\int_{{\bf K}^{\#}}f$ $\left(n^{\#}\left[\begin{smallmatrix}a\tau&0\\ 0&\tau^{-1}\end{smallmatrix}\right]\,k^{\#}\right)|a|_{\mathbb{A}}^{-2}|{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau)|_{\mathbb{A}}^{-2}$
@@
where ${\rm{d}}k^{\#}$ is the normalized Haar measure on ${\bf K}^{\#}=\prod_{p\leqslant\infty}{\bf K}_{p}^{\#}$. Note that the measure on $\mathsf{N}^{\#}({\mathbb{A}})\cong{\mathbb{A}}_{E}$ is $\eta_{{\mathbb{A}}_{E}}$ defined above, which is $\frac{\sqrt{|D|}}{2}\times\prod_{p}\eta_{E_{p}}\,(p\leqslant\infty)$ due to the formulas $(\prod_{p}\eta_{E_{p}})({\mathbb{A}}_{E}/E)=|D|^{1/2}/2$ ([49, Chap V §4 Proposition 7]) and $\eta_{{\mathbb{A}}_{E}}({\mathbb{A}}_{E}/E)=1$.

### v2

@@ 1. Introduction
Lately, the Bessel periods on the symplectic similitude group $\mathsf{G}:={\bf GSp}_{2}$ of rank $2$ progressively gain its importance in the arithmetic of Siegel modular forms. The notion of Bessel period and the associated local model of admissible representations of $\mathsf{G}$ were originally introduced by Novodvorski and Piatetski-Shaprio in their work ([27], [28], [31]), where a modern treatment of Andrianov’s integral representation ([1], [2]) of the spinor $L$-function for Siegel modular forms was developed. Related to this, we should also mention independent work by Sugano ([44]), which handles vector-valued Hilbert-Siegel cusp forms on arithmetic groups defined by maximal orders of indefinite division algebras. The Bessel period also plays an essential role in a formulation of Böcherer’s conjecture posed by Liu [24] in a style of the refined Gan-Gross-Prasad conjecture. This version of Böcherer’s conjecture has been completely solved by Furusawa and Morimoto ([10], [11]). In this paper, we compute the Rankin-Selberg type integral a la Piatetski-Shapiro for holomorphic vector-valued Siegel cusp forms of square-free levels explicitly, using results by Pitale-Schmidt [33] on local new vectors on $\mathsf{G}({\mathbb{Q}}_{p})\,(p<\infty)$; since we rely on the local-global method, our result is immediately applied to non-holomorphic Siegel cusp forms of discrete series type (cf. [21], [22]). We explain the moral of our investigation for scalar-valued Siegel modular forms of weight $l$. Our focus is an asymptotic formula as $N+l\rightarrow\infty$ of the average
@@
where $\mathsf{G}^{\#}$ is a subgroup of $\mathsf{G}$ defined by a quadratic field $E$ of discriminant $D<0$ (see §2.1), $E(s,\Lambda,\mu)\,(s\in{\mathbb{C}})$ is the Eisenstein series on $\mathsf{G}^{\#}({\mathbb{A}})$ attached to a ray class character $\Lambda$ of $E$ and a finite order character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}\rightarrow{\mathbb{C}}^{1}$ (see §3.1), $B^{\Lambda}(\varphi)$ is the Bessel period of $\varphi$ in an orthonormal basis ${\mathscr{B}}(l,N)$ of Siegel cusp forms of weight $l$ and square free level $\Gamma_{0}(N)$, and $\langle E(s,\Lambda,\mu)\mid\varphi\rangle$ is the $L^{2}$-inner product of $E(s,\Lambda,\mu)$ and $\varphi|\mathsf{G}^{\#}({\mathbb{A}})$. We suppose the ramifications of $E/{\mathbb{Q}}$, $\mu$, $\Lambda$, $\pi$ are mutually disjoint. Actually, for application ([20]), we consider a more general average than (1.1) by requiring ${\mathscr{B}}(l,N)$ to consist of joint eigenforms by Hecke operators and by putting the eigenvalue of such a Hecke operator at each $\varphi$ in the formation of the sum (1.1). With such a generality, we obtain a completely explicit expression of (1.1) in terms of
@@
  • An average of Fourier coefficient of a newform $\varphi_{\pi}^{0}$ in $\pi$, denoted by $R(\varphi_{\pi}^{0},E,\Lambda)$ (see (4.12)) introduced in [18].
@@
For the precise statement including the vector-valued case, we refer to Proposition 5.1 and Theorem 5.2. Owing to the explicit nature of our result, we are able to show that, asymptotically as the prime level $N$ grows to infinity with the weight $l$ being fixed, the contribution to (1.1) of the old forms, the Yoshida lifts and the Saito-Kurokawa lifts are 00 (see Theorem 5.4). These results are necessary in our forthcoming paper [20], where we shall work out the geometric side of a relative trace formula whose spectral side is essentially (1.1). We should remark that an average similar to our computed average (5.9) without the factors ${\widehat{L}}(s+1/2,\pi,\mu)$ and $\widehat{f_{S}}(\pi)$ is studied in [7, §3]. When $E={\mathbb{Q}}(i)$, our average (5.9) is almost the same as the one in [5, §8] except the appearance of the local periods.
@@
Let us explain the structure of this paper. In §2, we introduce basic objects related to algebraic subgroups of the symplectic group $\mathsf{G}$ and Haar measures on their $p$-adic, real and adelic points. In §3.1, the Rankin-Selberg type integral and the global Bessel function for Siegel cusp forms are recalled; after this preparation, the basic identity (Lemma 3.2) is stated, which computes the pairing $\langle E(s,\Lambda,\mu),\varphi\rangle$ in (1.1) in terms of the Mellin transform of the global Bessel function attached to $\varphi$. We review the proof ([31]) to determine a constant depending on Haar measures exactly. In §4, we extend the definition of the quantity $R(\varphi,E,\Lambda)$ for a ray class character $\Lambda$, which was originally considered in [18] for class group character; we describe its behavior under the Galois conjugation $\Lambda\mapsto\Lambda^{\dagger}$ (Lemma 4.6). The local multiplicity one theorem for Bessel models of $\pi_{p}$’s allows us to split $B^{\Lambda}(\varphi)$ for $\varphi\in\pi$ corresponding to a pure tensor in $\otimes_{p}\pi_{p}$ to an Euler product of local Bessel functions for $\pi_{p}$ up to a constant. To define our newform $\varphi_{\pi}^{0}$ and identify this constant with $R(\varphi_{\pi}^{0},E,\Lambda)$ (see Lemma 4.7), we fix pairs of a local Bessel functional $\ell_{p}$ on $\pi_{p}$ and a local new vector $\xi_{p}$ of $\pi_{p}$ such that $\ell_{p}(\xi_{p})=1$; in §4.4, we explain that the main result of [33] ensures the existence of such $\{(\ell_{p},\xi_{p})\}_{p}$ when $\pi$ corresponds to a new forms on $\Gamma_{0}(N)$. In §6, we compute the local zeta-integrals by using results of [33]. The unramified computations are known (see §6.2). Other cases, as well as the explicit determination of the local periods of prime level (see §6.9), seem to be new, so that some details are given in §6.6, §6.7, and §6.5. In §4.6, the functional equation of the $L$-function is deduced from the basic identity by using Lemma 4.6. Although a “nice” functional equation of the $L$-function itself is known in a broader generality ([41, Lemma 1.2]) owing to Arthur’s classification of the discrete spectrum, we include this section for the sake of completeness and to provide a proof free from Arthur’s result, confirming that the local $L$-factor and the local $\varepsilon$-factor defined by [31] coincides with the ones proposed in [38, Tables 2,3]. In section §5, we prove Proposition 5.1 and Theorem 5.2; by invoking a result by Furusawa-Morimoto ([10, Theorem 8.1]), we obtain an explicit expression of (1.1) solely in terms of $L$-functions (Theorem 5.3). To extend the result of Furusawa-Morimoto to a case when $\Lambda$ ramifies, we need a formula of a local period of the unramified matrix coefficient, we include some detail in §7 because this case is not handled in [7].
@@ 2.1. Embedding of groups
Let $E={\mathbb{Q}}(\sqrt{D})\subset{\mathbb{C}}$ be an imaginary quadratic field of discriminant $D<0$, so that $\sqrt{D}=i|D|^{1/2}$ with $i$ being the imaginary unit. Let $\mathfrak{o}_{E}$ denote the ring of integers of $E$. Choose a ${\mathbb{Z}}$-basis $\{1,\theta\}$ of $\mathfrak{o}_{E}$, so that
@@
The different ${\mathfrak{d}}_{E/{\mathbb{Q}}}$, defined as ${\mathfrak{d}}_{E/{\mathbb{Q}}}^{-1}:=\{\tau\in E\mid{\operatorname{tr}}_{E/{\mathbb{Q}}}(\tau\mathfrak{o}_{E})\subset{\mathbb{Z}}\}$, is ${\sqrt{D}}\mathfrak{o}_{E}$; thus a symplectic ${\mathbb{Z}}$-bi-linear form $\langle\cdot,\cdot\rangle:\mathfrak{o}_{E}^{2}\times\mathfrak{o}_{E}^{2}\longrightarrow{\mathbb{Z}}$ is defined by
@@
Then $\iota_{\theta}(h)\in\mathsf{G}(R)$ and $\nu(\iota_{\theta}(h))=\det h$. Thus, we have an embedding
@@
Let ${\mathsf{B}}^{\#}$ be the Borel subgroup of $\mathsf{G}^{\#}$ such that ${\mathsf{B}}^{\#}(R)$ coincides with the set of all points
@@
Define a bilinear form on ${\mathsf{V}}(R)$ as $(X,Y):=-{\operatorname{tr}}(XY^{\dagger})$ and ${\bf SO}({\mathsf{V}}(R))$ as the special orthogonal group of (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ).
In the proof of Lemma 4.3, we use the following lemma:
Lemma 2.1.
Let $F$ be a field with characteristic 00.
  • (i)
    The map ${\mathsf{s}}:{\bf GL}_{2}(F)\to{\bf GL}({\mathsf{V}}(F))$ induces an isomorphism
    ${\bf PGL}_{2}(F)\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}{\bf SO}({\mathsf{V}}(F)).$
  • (ii)
    The map ${\mathsf{s}}\circ I_{\theta}:(E\otimes_{\mathbb{Q}}F)^{\times}/F^{\times}\to{\rm Stab}_{{\bf SO}({\mathsf{V}}(F))}(T_{\theta})$ is an isomorphism.
@@ 2.2. Open compact subgroups at finite places
Let $p$ be a prime number. Set $E_{p}:=E\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ and $\mathfrak{o}_{E,p}:=\mathfrak{o}_{E}\otimes_{\mathbb{Z}}{\mathbb{Z}}_{p}$. Then, ${\bf K}^{\#}_{p}:=\mathsf{G}^{\#}({\mathbb{Z}}_{p})$, which coincides with the stabilizer in $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ of the $\mathfrak{o}_{E,p}$-lattice $\mathfrak{o}_{E,p}^{2}\subset E_{p}^{2}$, is a maximal compact subgroup of $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$, and ${\bf K}_{p}:=\mathsf{G}({\mathbb{Z}}_{p})$ is the standard maximal compact subgroup of $\mathsf{G}({\mathbb{Q}}_{p})$. For a non-zero ideal ${\mathfrak{n}}\subset{\mathbb{Z}}_{p}$, set
@@
For a non-zero ideal ${\mathfrak{n}}\subset{\mathbb{Z}}_{p}$, we have ${\bf K}_{0}^{\#}({\mathfrak{n}}\mathfrak{o}_{E,p})=\iota_{\theta}^{-1}({\bf K}_{0}({\mathfrak{n}}))$.
@@ 2.3. Maximal compact subgroup at the archimedean place
Note that ${\bf K}_{\infty}\subset{\bf Sp}_{2}({\mathbb{R}})$. The action of $\mathsf{G}({\mathbb{R}})^{0}$ on the Siegel upper-half space ${\mathfrak{h}}_{2}:=\{Z\in{\bf M}_{2}({\mathbb{C}})\mid{}^{t}Z=Z,\,{\operatorname{Im}}(Z)\gg 0\}$, denoted by $(g,Z)\mapsto g\langle Z\rangle$ is defined by the usual formula, i.e.,
@@ 2.4. Haar measures
For locally compact unimodular topological groups $H$ relevant to us, we fix Haar measures $\eta_{H}$ on $H$ in the following manner. Let ${\mathbb{A}}$ be the adele ring of ${\mathbb{Q}}$ and $\psi:{\mathbb{A}}/{\mathbb{Q}}\rightarrow{\mathbb{C}}^{1}$ the basic character.
@@
Let $p\leqslant\infty$. Any element $g^{\#}\in\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ is written as
(2.19) $g^{\#}=\left[\begin{smallmatrix}1&\beta\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}\tau&0\\ 0&\bar{\tau}a\end{smallmatrix}\right]\,k^{\#}$
with $a\in{\mathbb{Q}}_{p}^{\times}$, $\tau\in E_{p}^{\times}$, $\beta\in E_{p}$ and $k^{\#}\in{\bf K}_{p}^{\#}$. Then, our Haar measure on $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ is symbolically defined as
(2.20) ${\rm{d}}\eta_{\mathsf{G}^{\#}({\mathbb{Q}}_{p})}(g_{p}^{\#})=|a|_{p}^{2}\times{\rm{d}}\eta_{E_{p}}(\beta)\,{\rm{d}}\eta_{{\mathbb{Q}}_{p}^{\times}}(a)\,{\rm{d}}\eta_{E_{p}^{\times}}(\tau)\,{\rm{d}}\eta_{{\bf K}_{p}^{\#}}(k^{\#}),$
where ${\rm{d}}\eta_{{\bf K}_{p}^{\#}}(k^{\#})$ is the Haar measure on ${\bf K}_{p}^{\#}$ with volume $1$. For $p<\infty$, the Haar measure $\eta_{\mathsf{G}^{\#}({\mathbb{Q}}_{p})}$ is the one with ${\operatorname{vol}}({\bf K}_{p}^{\#})=1$. On the adele group $\mathsf{G}^{\#}({\mathbb{A}})$, we use the product measure of Haar measures on $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ for $p\leqslant\infty$. Then, from (2.20), we get
(2.21) $\int_{\mathsf{G}^{\#}({\mathbb{A}})}f(g^{\#})\,{\rm{d}}g^{\#}=\tfrac{\sqrt{|D|}}{2}\int_{\mathsf{N}^{\#}({\mathbb{A}})}\int_{{\mathbb{A}}^{\times}}\int_{{\mathbb{A}}^{\times}_{E}}\int_{{\bf K}^{\#}}f$ $\left(n^{\#}\left[\begin{smallmatrix}\tau&0\\ 0&a\bar{\tau}\end{smallmatrix}\right]\,k^{\#}\right)|a|_{\mathbb{A}}^{2}$
@@
where ${\rm{d}}k^{\#}$ is the normalized Haar measure on ${\bf K}^{\#}=\prod_{p\leqslant\infty}{\bf K}_{p}^{\#}$. Note that the measure on $\mathsf{N}^{\#}({\mathbb{A}})\cong{\mathbb{A}}_{E}$ is $\eta_{{\mathbb{A}}_{E}}$ defined above, which is $\frac{\sqrt{|D|}}{2}\times\prod_{p}\eta_{E_{p}}\,(p\leqslant\infty)$ due to the formulas $(\prod_{p}\eta_{E_{p}})({\mathbb{A}}_{E}/E)=|D|^{1/2}/2$ ([50, Chap V §4 Proposition 7]) and $\eta_{{\mathbb{A}}_{E}}({\mathbb{A}}_{E}/E)=1$.